OpenStudy (jacky90):

In an electrical circuit, the current passing through a conductor varies inversely with the resistance. Suppose that when the current is 3 A (amperes), the resistance is 40ohms. What is the current when the resistance is 15 ohms?

OpenStudy (sooobored):

Ohm's law \[V=IR\] where V is voltage in volts where I is current in amps where R is resistance in ohms

OpenStudy (danjs):

Recall an inverse relationship , Y is inversely proportional to X \[\large y~~~\alpha~~~\frac{ 1 }{ x }\] insert a constant value k to make an equation, current y varies inversely with the resistance x. \[\large y=\frac{ k }{ x }\]

OpenStudy (danjs):

The data point they give you allows you to figure out the value of the constant k.. If current Y=3, then the resistance is x=40 3=k/40 k=120 so the equation is \[\large y=\frac{ 120 }{ x }\]

OpenStudy (danjs):

for the resistance is instead now x=15 ohms, you can calculate the current

OpenStudy (jacky90):

Im getting 8 amperes? is that right?

OpenStudy (danjs):

yes

OpenStudy (jacky90):

TY :)